# MolTailor: Tailoring Chemical Molecular Representation to Specific Tasks via Text Prompts

### v1

@@ Abstract
Deep learning is now widely used in drug discovery, providing significant acceleration and cost reduction. As the most fundamental building block, molecular representation is essential for predicting molecular properties to enable various downstream applications. Most existing methods attempt to incorporate more information to learn better representations. However, not all features are equally important for a specific task. Ignoring this would potentially compromise the training efficiency and predictive accuracy. To address this issue, we propose a novel approach, which treats language models as an agent and molecular pretraining models as a knowledge base. The agent accentuates task-relevant features in the molecular representation by understanding the natural language description of the task, just as a tailor customizes clothes for clients. Thus, we call this approach MolTailor. Evaluations demonstrate MolTailor’s superior performance over baselines, validating the efficacy of enhancing relevance for molecular representation learning. This illustrates the potential of language model guided optimization to better exploit and unleash the capabilities of existing powerful molecular representation methods. Our codes and appendix are available at https://github.com/SCIR-HI/MolTailor.
@@ Introduction
To achieve this goal, we adopt a dual-tower model structure with one language pretraining module and one molecular pretraining module, joint together through a cross-attention module. Meanwhile, we constructed a new pretraining task: Molecule-Text Multi-Task Regression (MT-MTR). The dataset for this task consists of (molecule, task description, regression labels) triplets. Here, the molecule is represented as a SMILES string (Weininger 1988), the task description is a text prompt describing molecular properties that are most helpful for solving the task, and the regression labels are values of features mentioned in the task description. The model needs to predict the regression labels based on the SMILES and task description. This pretraining task teaches the model to enhance the weights of task-relevant features in the molecular representation according to the task description. Our contributions can be summarized as follows:
@@
Molecular pretraining models
@@
Step 2: Calculate molecular properties. For each molecule obtained in the previous step, we use RDKit to calculate 209 properties [^2] RDKit’s Descriptor module provides 209 descriptors, which we uniformly refer to as “properties” in this paper for clarity.. Among these properties, some are continuous values while others are discrete values. We unify them into regression tasks.
@@
Table 1: Prompt template for interacting with GPT to generate task descriptions
@@
where $f^{r}$ denotes reshape function, [*]delimited-[][*][ * ] denotes concatenation along the dimension corresponding to the sequence length $n$. The output of the MT-Encoder Block is computed:
@@
When applying MolTailor to downstream tasks, we use GPT-4 to analyze the specific task and then generate the corresponding task description, as shown in Fig. 2e. We use the template in Tab. 1 to prompt GPT to generate descriptions that mimic the format of those in the pretraining corpus. At the same time, we restrict it to select task-relevant features from the set of properties supported by RDKit. Then, we feed the generated analysis and SMILES into MolTailor to obtain the task-specific molecular representation:
@@
Q1: Are the task-specific molecular representations generated by MolTailor better than general representations? Q2: Can MolTailor achieve performance improvements on different M-Encoders? Q3: How do the task descriptions influence the effectiveness of pretraining? Q4: How do different text prompts affect MolTailor? Q5: When MolTailor achieves better performance, are the task-relevant features in the molecular representations enhanced? Q6: Does MolTailor pay attention to the key information in both the molecules and text prompts?
@@
Table 2: Evaluation results of MolTailor and baselines under the linear probe setting on MoleculeNet. Here, ME stands for Molecule Encoder and TE stands for Text Encoder in MoMu-ME/TE. $*{}^{*}$ denotes using ChemBERTa-2 as the M-Encoder, while MolTailor indicates using CHEM-BERT. The table shows the average performance over 3 runs. Standard deviations are omitted due to space limitations but can be viewed in the appendix. Results in bold indicate the best performance. Green highlights cases where MolTailor outperforms the backbone, while yellow indicates the opposite.
@@
Q1: We evaluate whether molecular representations enhanced by task descriptions could improve performance across 8 tasks. Tab. 2 shows that MolTailor achieves performance gains over the backbone model on the 4 regression tasks, and notably attains state-of-the-art (SOTA) results on ESOL, FreeSolv, and Lipophilicity datasets. On the 4 classification tasks, MolTailor’s performance is inconsistent, with gains on HIV and Tox21 but losses on the other 2 datasets when using CHEM-BERT as backbone. We hypothesize this may be because the pretraining task favors regression, which could inhibit performance on classification. And this implies classification and regression tasks may require different optimizing approaches. Additionally, we test converting MT-MTR to MT-MTC for pretraining but see minimal improvement, which indicates this is not due to the form of the task.
@@
Alation Study (Q3 & Q4 & Q5)
Q3. To evaluate the influence of task descriptions on pretraining in MT-MTR, we first remove the task descriptions to obtain the dataset $*{}^{*}$, then pretrain CHEM-BERT and PubMedBERT on $*{}^{*}$ to get $*{}^{*}$ and $*{}^{*}$ models. Next, we concatenate the task descriptions after the SMILES strings to construct dataset ${}^{\dagger}$, and pretrain PubMedBERT on ${}^{\dagger}$ to obtain ${}^{\dagger}$. As shown in Tab. 3, $*{}^{*}$ improves over untrained CHEM-BERT on regression tasks but drops on classification, indicating MTR as a pretraining task benefits downstream regression but negatively impacts classification. Also, ${}^{\dagger}$ over $*{}^{*}$ and MolTailor over CHEM-BERT both demonstrate further regression performance gains but classification performance drops. This shows introducing task descriptions helps models better learn from the data. That is, if constructing new labels can improve model performance on classification tasks, then supplementing task descriptions can further amplify such gains.
Models Classificationnormal-$\uparrow$ Regressionnormal-$\downarrow$
@@
Table 3: Experimental results for Q3. For these experiments, the overlapping data are excluded, and all models are pretrained for only 20 epochs.
Q4. We replace the task-specific descriptions generated via GPT-4 analysis with the irrelevant text “to be or not to be, this is the question.” as noise prompts. Results in Tab. 4 show degraded performance with noise prompts, indicating information in the task descriptions does help the model generate better molecular representations. However, it should be noted that the performance drop is slight, suggesting textual information is not as important as expected in the process of generating representations. Further investigation into this phenomenon is needed.
Prompt Types Classificationnormal-$\uparrow$ Regressionnormal-$\downarrow$
@@
Q5. Earlier experiments have shown that task-specific molecular representations outperform general ones. Now, we aim to test if task-relevant features are more pronounced in these task-specific representations. Our theory is simple: if a feature is more prominent in the representation, it should lead to better predictions for that feature compared to a less distinct representation.
For our experiment, we chose the ESOL dataset and four models: KV-PLM, CHEM-BERT, MolTailor, and MolTailor (noise), where the latter refers to a version using a noise prompt. We calculated 209 properties for each molecule in ESOL, creating a new dataset named ESOL-MTR, comprising pairs of SMILES strings and their 209 regression labels. We trained all four models on ESOL-MTR and evaluated their ability to predict the 209 properties of the ESOL molecules. From these, we focused on 4 properties for Fig. 3. We calculated the Normalized RMSE by subtracting the average RMSE of the four models on a task from the original RMSE and then dividing it by the standard deviation. The four properties we selected include two mentioned in the task description (MolWt and FractionCSP3), one relevant but unmentioned (EState_VSA3), and one unrelated and unmentioned (Kappa1). Our analysis led to two main observations:
  • For the properties mentioned in the task description, MolTailor outperforms MolTailor (noise), confirming that these mentioned properties are indeed emphasized.
  • We find prompts influence both mentioned and unmentioned properties. “EState_VSA3” is a relevant but unmentioned property shows performance gains. In contrast, “Kappa1” is an irrelevant and unmentioned property performance decline. The results show that by understanding text prompts, the model can attend to properties not explicitly stated.
Figure 3: Normalized RMSE of four models on ESOL task and associated/unassociated proterties. The x-axis shows the names of tasks including ESOL and four properties, of which the first three are related to ESOL, while the last is not. The y-axis plots N-RMSE for the four models on each task, with lower values indicating better performance.
@@
Q6. We analyze the attention matrices from the last layers of the UT-Encoder and MT-Encoder to investigate whether MolTailor pays attention to key information. The analysis results are shown in Fig. 4. Specifically, we select ESOL as a representative task related to solubility. If the model attends to information like molecular weight and polar functional groups critical for determining solubility, it suggests key information is captured.
After training on ESOL, we input SMILES from the test set and task description into the model to obtain attention matrices $\boldsymbol{M}_{UT}$ and $\boldsymbol{M}_{MT}$. Since we use the vector corresponding to the “[CLS]” token for prediction, we believe the larger the attention value between “[CLS]” and a token in $\boldsymbol{M}_{UT}$ or $\boldsymbol{M}_{MT}$, the more the model focuses on that information.
In Fig. 4, tokens highlighted in red in the top 5 most attended by the UT/MT-Encoders relate to molecular weight or polar functional groups, indicating the model does pay attention to key textual information. Also, the two molecular graphs plotted using the most attended molecular tokens by the MT-Encoder under different prompts show the original prompt leads to more attention on polar functional groups like ester and ketone, meaning critical information is better captured.
@@
Overall, this work provides a new perspective on molecular representation learning, not only striving to include more comprehensive information in the representations, but also combining contextual information and obtaining molecular representations more suitable for specific tasks through tailoring. In this work, we not only utilize the knowledge implied in the text modality, more importantly, we try to leverage the excellent reasoning ability of language models, which will become more important in the age of large models.

### v2

@@ Abstract
Deep learning is now widely used in drug discovery, providing significant acceleration and cost reduction. As the most fundamental building block, molecular representation is essential for predicting molecular properties to enable various downstream applications. Most existing methods attempt to incorporate more information to learn better representations. However, not all features are equally important for a specific task. Ignoring this would potentially compromise the training efficiency and predictive accuracy. To address this issue, we propose a novel approach, which treats language models as an agent and molecular pretraining models as a knowledge base. The agent accentuates task-relevant features in the molecular representation by understanding the natural language description of the task, just as a tailor customizes clothes for clients. Thus, we call this approach MolTailor. Evaluations demonstrate MolTailor’s superior performance over baselines, validating the efficacy of enhancing relevance for molecular representation learning. This illustrates the potential of language model guided optimization to better exploit and unleash the capabilities of existing powerful molecular representation methods. Our code is available at https://github.com/SCIR-HI/MolTailor.
@@ Introduction
To achieve this goal, we adopt a dual-tower model structure with one language pretraining module and one molecular pretraining module, joint together through a cross-attention module. Meanwhile, we constructed a new pretraining task: Molecule-Text Multi-Task Regression (MT-MTR). The dataset for this task consists of (molecule, task description, regression labels) triplets. Here, the molecule is represented as a SMILES string (Weininger 1988), the task description is a text prompt describing molecular properties that are most helpful for solving the task, and the regression labels are values of properties mentioned in the task description. The model needs to predict the regression labels based on the SMILES and task description. This pretraining task teaches the model to enhance the weights of task-relevant properties in the molecular representation according to the task description. Our contributions can be summarized as follows:
@@
Molecular Pretraining Models.
@@
Step 2: Calculate molecular properties. For each molecule obtained in the previous step, we use RDKit to calculate 209 properties [^2] RDKit’s “Descriptor” module provides 209 descriptors, which we uniformly refer to as “properties” in this paper for clarity.. Among these properties, some are continuous values while others are discrete values. We unify them into regression tasks.
@@
Table 1: Prompt templates for interacting with GPT to generate task descriptions.
@@
where $f^{r}$ denotes reshape function, []delimited-[][*][ ∗ ] denotes concatenation along the dimension corresponding to the sequence length $n$. The output of the MT-Encoder Block is computed:
@@
When applying MolTailor to downstream tasks, we use GPT-4 to analyze the specific task and then generate the corresponding task description, as shown in Fig. 2e. We use the template in Tab. 1 to prompt GPT to generate descriptions that mimic the format of those in the pretraining corpus. Notably, we restrict it such that it can only select the names of properties from the set supported by RDKit. Then, we feed the generated analysis and SMILES into MolTailor to obtain the task-specific molecular representation:
@@
Q1: Are the task-specific molecular representations generated by MolTailor better than general representations? Q2: Can MolTailor achieve performance improvements on different M-Encoders? Q3: How do the task descriptions influence the effectiveness of pretraining? Q4: How do different text prompts affect MolTailor? Q5: When MolTailor achieves better performance, are the task-relevant properties in the molecular representations enhanced? Q6: Does MolTailor pay attention to the key information in both the molecules and text prompts?
@@
Table 2: Evaluation results of MolTailor and baselines under the linear probe setting on MoleculeNet. Here, ME stands for Molecule Encoder and TE stands for Text Encoder in MoMu-ME/TE. MolTailor denotes using ChemBERTa-2 as the M-Encoder, while MolTailor indicates using CHEM-BERT. The table shows the average performance over 3 runs. Standard deviations are omitted due to space limitations but can be viewed in the appendix. Results in bold indicate state-of-the-art (SOTA) performance, while the underlined results show that the model outperforms the backbone used in the M-Encoder.
@@
Q1: We evaluate whether molecular representations enhanced by task descriptions could improve performance across 8 tasks. Tab. 2 shows that MolTailor achieves performance gains over the backbone model on the 4 regression tasks, and notably attains state-of-the-art (SOTA) results on ESOL, FreeSolv, and Lipophilicity datasets. On the 4 classification tasks, MolTailor’s performance is inconsistent, with gains on HIV and Tox21 but losses on the other 2 datasets when using CHEM-BERT as backbone. We hypothesize that the difference in model performance on classification and regression tasks is due to the pretraining data benefiting the regression task more. Meanwhile, we experimented with converting the regression pretraining task into a classification form. The results were still similar, eliminating the influence of the form of the pretraining task on downstream tasks.
@@
Ablation Study (Q3 & Q4 & Q5)
Q3. To evaluate the influence of task descriptions on pretraining in MT-MTR, we first remove the task descriptions to obtain the dataset MT-MTR, then pretrain CHEM-BERT and PubMedBERT on MT-MTR to get CHEM-BERT and PubMedBERT models. Next, we concatenate the task descriptions after the SMILES strings to construct dataset MT-MTR, and pretrain PubMedBERT on MT-MTR to obtain PubMedBERT. As shown in Tab. 3, CHEM-BERT improves over untrained CHEM-BERT on regression tasks but drops on classification, indicating MTR as a pretraining task benefits downstream regression but negatively impacts classification. Also, PubMedBERT over PubMedBERT and MolTailor over CHEM-BERT both demonstrate further regression performance gains but classification performance drops. This shows introducing task descriptions helps models better learn from the data. That is, if constructing new labels can improve model performance on classification tasks, then supplementing task descriptions can further amplify such gains.
Q4. We replace the task-specific descriptions generated via GPT-4 with the irrelevant text “to be or not to be, this is the question.” as noise prompts. Results in Tab. 4 show degraded performance with noise prompts, indicating information in the task descriptions does help the model generate better molecular representations. However, it should be noted that the performance drop is slight, suggesting textual information is not as important as expected in the process of generating representations. Further investigation into this phenomenon is needed.
[…]
@@
Table 3: Experimental results for Q3. CHEM-BERT and PubMedBERT use their original weights when evaluated on downstream tasks. The remaining methods are pretrained on MT-MTR and its variants, with overlapping molecules between pretraining and downstream data removed. Additionally, all methods are pretrained for only 20 epochs.
[…]
@@
Figure 3: Performance of three methods on ESOL and molecular properties prediction tasks for Q5. The x-axis shows the task names, the y-axis shows the normalized RMSE, with lower values indicating better performance. Of the four molecular properties, the first three are related to the ESOL task, while the last one is opposite.
Q5. The experiments in Tab. 2 show that the molecule representations generated by MolTailor attained better performance. However, are the obtained representations more task-specific, meaning the task-relevant properties in the representations are enhanced? To validate this, we designed the following experiment: we used the representations generated by different methods to predict task-related and task-unrelated properties. If the expectations are met, then the representations enhanced by MolTailor should have better performance in predicting task-related molecular properties.
We conduct experiments on the ESOL dataset. For the models, we use MolTailor and its corresponding backbone CHEM-BERT. We also test the performance of MolTailor using irrelevant text as the prompt, denoted as MolTailor(noise). In detail, we first use RDKit to compute the properties of the molecules in the ESOL dataset, obtaining the ESOL-MTR dataset. Then, we randomly split the dataset into 8:1:1 for training, validation, and testing.
The results are shown in Fig. 3. We present the performance of the three methods on the ESOL task and in predicting four molecular properties. Among the four selected properties, two have names mentioned in the prompt and are relevant to the ESOL task (MolWt and FractionCSP3), one is not mentioned but is relevant (EState_VSA3), and one is neither mentioned nor relevant (Kappa1). We use RMSE as the evaluation metric. For better visualization, we normalize the RMSE under each task using the formula: Normalized $\text{Normalized RMSE}=\tfrac{\text{RMSE}-\text{mean}(\text{RMSE})}{\text{std}(\text{RMSE})}$.
The results show that: a) Compared to CHEM-BERT, MolTailor does focus more on the properties mentioned in the prompt, validating that the obtained representations are indeed task-specific. b) MolTailor not only pays more attention to the relevant properties explicitly stated in the prompt, but also to the unstated yet still relevant properties. Meanwhile, it decreases attention on the unrelated, unmentioned properties. This demonstrates the method’s generalization capability. c) The declined performance of MolTailor(noise) in property prediction indicates that the prompt does help the model attend to critical information.
@@
Q6. We analyze the attention matrices from the last layers of the UT/MT-Encoder to investigate whether MolTailor pays attention to key information. We select the ESOL dataset, which is related to solubility, to conduct experiments. If the model attends to information such as molecular weight or polar functional groups that are critical determinants of solubility, it suggests key information is captured.
In detail, We pass the SMILES from the test set through MolTailor trained on ESOL to obtain the required attention matrices $\boldsymbol{M}_{UT}$ and $\boldsymbol{M}_{MT}$. Then, we take the values of the matrix between the “[CLS]” token and the other tokens as the required attention weights.
The analysis results in Fig. 4 show that MolTailor does pay attention to solubility-related information, such as “groups”, “ethers”, and other highlighted tokens. Notably, the tokens here are merged to form complete words. Additionally, the model also pays attention, as seen in the left molecule attention graph, to polar functional groups within the SMILES. Furthermore, variations between the two molecule attention graphs under different prompts indicate effective prompts help the model better focus on key information that assist in solving the task
@@
Overall, in this work we propose a new perspective on molecular representation learning. Instead of trying to incorporate more information into the representations, we obtain more task-specific representations by combining contextual information. At the same time, we not only utilize the knowledge contained in the text modality, but also try to leverage the reasoning ability of language models, which has huge potential in the era of large language models.